the perimeter of a rectangular garden is 64 feet. if the length is 4 feet less than twice its width, what are the dimensions of the garden?
Answer by sachi(548)

let the width = x
then length = 2x-4
the perimeter = 2(x+2x-4)
so 2(x+2x-4)=64
or 3x-4=64/2=32
or 3x=32+4=36
or x=36/3=12=width
length=2*12-4=20
ans